the probability of obtaining even prime number on each die,when a pair of rice is rolled is

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Total number of ways =6*6=36
2 is the only even prime number
The probability is given by =(1/6)(1/6)=1/36
Thanks and Regards
Shaik Aasif
askIITians faculty